# Upper and Lower Bounds for the Linear Ordering Principle

### v1

@@ Abstract
Korten and Pitassi (FOCS, 2024) defined a new [^1] Note that this notation had been used in the past [Sch83] for a very different class, which has been apparently forgotten after that. complexity class $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ as the polynomial-time Turing closure of the Linear Ordering Principle. They asked whether a Karp–Lipton–style collapse can be proven for $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$. We answer this question affirmatively by showing that $\mathbf{{P}}^{\mathbf{{prMA}}}\subseteq\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$. As a byproduct, we also answer an open question of Chakaravarthy and Roy (Computational Complexity, 2011) whether $\mathbf{{P}}^{\mathbf{{prMA}}}\subseteq\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$.
We complement this result by providing a new upper bound for $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$, namely $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}\subseteq\mathbf{{P}}^{\mathbf{{prSBP}}}$. Thus we are placing $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ between $\mathbf{{P}}^{\mathbf{{prMA}}}$ and $\mathbf{{P}}^{\mathbf{{prSBP}}}$. One technical ingredient of this result is an algorithm that approximates the number of satisfying assignments of a Boolean circuit using a $\mathbf{{prSBP}}$ oracle (i.e. in $\mathbf{{FP}}^{\mathbf{{prSBP}}}$), which could be of independent interest.
Finally, we prove that $\mathbf{{P}}^{\mathbf{{prO_{2}^{\mathbf{{\scriptscriptstyle P}}}}}}\subseteq\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$, which implies that the Karp–Lipton–style collapse to $\mathbf{{P}}^{\mathbf{{prOMA}}}$ is actually better than both known collapses to $\mathbf{{P}}^{\mathbf{{prMA}}}$ due to Chakaravarthy and Roy (Computational Complexity, 2011) and to $\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ also due to Chakaravarthy and Roy (STACS, 2006).
@@ 1 Introduction
The seminal theorem of Richard M. Karp and Richard J. Lipton [KL80] connected non-uniform and uniform complexity by demonstrating a collapse of the Polynomial Hierarchy assuming $\mathbf{{NP}}$ has polynomial-size Boolean circuits. This collapse has since been very instrumental in transferring lower bounds against Boolean circuits of fixed-polynomial [^2] That is, for any $k\in\mathbb{N}$, the class contains a language that cannot be computed by Boolean circuits of size $n^{k}$, i.e. a language outside of $\mathsf{Size}[n^{k}]$. size to smaller classes of the Polynomial Hierarchy. Since then, these results were strengthened in many ways yielding “minimal” classes that have such lower bounds and to which the Polynomial Hierarchy collapses.
@@ 1.1.1 Classes Based on Symmetric Alternation
An important notion in this context is that of symmetric alternation. Namely, one of the best collapses was based on the following idea ([Cai07], attributed to Sengupta): if polynomial-size circuits for SAT exist, two provers (defending the answers ‘yes’ and ‘no’, respectively) send such circuits to a polynomial-time bounded verifier who can use them to verify membership in any language in $\mathbf{{PH}}$. The corresponding class $\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ [Can96, RS98] was thus shown to have fixed-polynomial circuit lower bounds.
Indeed, since $\mathbf{{NP}}\subseteq\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$, if SAT requires super-polynomial circuits, we are done. Otherwise, the Polynomial Hierarchy, which is known to contain “hard” languages [^3] For every $k\in\mathbb{N}:\mathbf{{PH}}\not\subseteq\mathsf{Size}[n^{k}]$. by Kannan’s theorem [Kan82], collapses to $\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ and so do these hard languages. This technique has been known as a win-win argument in the literature [Kan82, BCG+96, KW98, Vin05, Cai07, San09, CR11, Vol14, IKV23].
@@
An important feature of these new results was that they were based on reducing finding a hard function to a total search problem. Namely, the works of Korten [Kor21] and Li [Li24] reduced the question to the so-called Range Avoidance problem. That is, given a function $f\colon\{0,1\}^{n}\to\{0,1\}^{m}$ with $m>n$, represented by a Boolean circuit, find a point outside its image. In [CHR24, Li24], the range avoidance problem has been reduced to symmetric alternation. Subsequently, Korten and Pitassi [KP24] reduced the range avoidance problem to the Linear Ordering Principle: given an implicitly described ordering relation, either find the smallest element or report a breach of the linear order axioms. A polynomial-time Turing closure of this principle gave rise to a new class $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}\subseteq\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$: a version of $\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ where the two provers provide points of a polynomial-time verifiable linear order on binary strings of a certain length (each point starting with the corresponding answer 00 or $1$), and the prover that provides the smaller element wins.
@@ 1.1.2 Classes Based on Merlin-Arthur Protocols
In a parallel line of research, the same questions were considered for classes based on Merlin–Arthur proofs: Santhanam [San09] has shown fixed-polynomial lower bounds for promise problems possessing such proofs (i.e. the class $\mathbf{{prMA}}$). In [CR11], Chakaravarthy and Roy have shown a Karp–Lipton–style collapse and thus fixed-polynomial size lower bounds for the class $\mathbf{{P}}^{\mathbf{{prMA}}}$. In particular, they presented a new upper bound for $\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ by showing that $\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}\subseteq\mathbf{{P}}^{\mathbf{{prAM}}}$. Combining this with a result of [AKSS95], that $\mathbf{{NP}}\subseteq\mathbf{{P/poly}}$ implies an “internal collapse” $\mathbf{{MA}}=\mathbf{{AM}}$ [^4] The internal collapse goes through for the promise versions of the classes as well., they concluded that the Polynomial Hierarchy collapses all the way to $\mathbf{{P}}^{\mathbf{{prMA}}}$. Subsequently, by applying the win-win argument, they obtained fixed-polynomial bounds for $\mathbf{{P}}^{\mathbf{{prMA}}}$, which (unlike $\mathbf{{prMA}}$) is a class of languages. It is to be noted though that since $\mathbf{{prMA}}$ is not a class of languages — while $\mathbf{{P}}^{\mathbf{{prMA}}}$ is, there is no immediate way to carry any lower bound against $\mathbf{{prMA}}$ over to $\mathbf{{P}}^{\mathbf{{prMA}}}$. Nonetheless, the relationship between $\mathbf{{P}}^{\mathbf{{prMA}}}$ and the classes of symmetric alternation (including $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$, $\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$, and $\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$) remained unknown.
1.1.3 $\mathbf{{SBP}}$, Approximate Counting, and Set Size Estimation
There is a class that sits between the two basic classes based on Arthur–Merlin protocols, yet its definition is not based on these protocols. The class $\mathbf{{SBP}}$, which stands for small bounded-error polynomial-time computations, was introduced in [BGM06] by Böhler et al. as a relaxation of the class $\mathbf{{BPP}}$ to the case when the acceptance probability is not required to be bounded away from 00. Indeed, this relaxation gives more power and as a result, as was shown in [BGM06], $\mathbf{{MA}}\subseteq\mathbf{{SBP}}\subseteq\mathbf{{AM}}$. Furthermore, $\mathbf{{SBP}}$ is the only known natural class that lies between $\mathbf{{MA}}$ and $\mathbf{{AM}}$.
Computing the number of accepting paths of a given non-deterministic Turing machine is a fundamental problem captured by the “counting” class #$\mathbf{\#\mathbf{{P}}}$. Yet, this class appears to be too powerful since, by Toda’s Theorem [Tod91], even a single query to it suffices to decide any language in the polynomial hierarchy, $\mathbf{{PH}}\subseteq\mathbf{{P}}^{\mathbf{\#\mathbf{{P}}[1]}}$! Given that, it is natural to explore approximations. To this end, one can consider the problem of Approximate Counting which refers to the task of approximating the number of accepting paths (within a constant factor). Equivalently, this problem can be framed as approximating the size of a set $S$ represented as the set of satisfying assignments of a Boolean circuit $C$. Previously, it was shown that this task could be carried out by a randomized algorithm using an $\mathbf{{NP}}$ oracle ($\mathbf{{FBPP}}^{\mathbf{{NP}}}$) [Sto85, JVV86] and by a deterministic algorithm using a $\mathbf{{prAM}}$ oracle ($\mathbf{{FP}}^{\mathbf{{prAM}}}$) [Sip83, GS86]. Shaltiel and Umans [SU06] show how to accomplish this task in $\mathbf{{FP}}^{\mathbf{{NP}}}$, yet under a derandomization assumption. All of these algorithms can be implemented using parallel oracle queries.
The decision version of the problem is to choose between two constant-factor estimates of the set size. For concreteness, consider the following problem called set-size estimation (or SSE for short): Given a set $S$ (via a Boolean circuit $C$) and an integer $m$ with a promise that either $\left|{S}\right|\geq m$ or $\left|{S}\right|\leq m/2$, our goal is to decide which case holds. Interestingly, a result of [BGM06] implies that this problem is complete for the corresponding class $\mathbf{{prSBP}}$ of promise problems (and thus it is hard for $\mathbf{{SBP}}$)! In terms of upper bounds, in a seminal paper [GS86], Goldwasser and Sipser have exhibited an Arthur–Merlin protocol not only for this problem, but also for the case when the set $S$ is represented by a non-deterministic circuit [^5] A non-deterministic circuit $C(x,w)$ accepts $x$ if there exists a witness $w$ for which $C(x,w)=1$.! This more general version of the problem (WSSE, where the set $S$ is given via a non-deterministic circuit), is complete for the class $\mathbf{{prAM}}$. (See Definition 5.2 for the formal definition of the problems; in fact, the factor-of-two gap in the estimates is arbitrary and can be replaced by any positive constant. .) In fact, Goldwasser–Sipser’s protocol proves the containment both for languages ($\mathbf{{SBP}}\subseteq\mathbf{{AM}}$) and for promise problems ($\mathbf{{prSBP}}\subseteq\mathbf{{prAM}}$). At the same time, it is important to highlight the distinction between the two versions of the problem — i.e. for the “standard” (SSE) vs non-deterministic (WSSE) Boolean circuits — which appears to be (at the very least) non-trivial. Notably, the work of [BGM06] established an oracle separation between $\mathbf{{SBP}}$ and $\mathbf{{AM}}$.
@@ 1.2 Promise Problems as Oracles
An important note is due on the use of a promise problem as an oracle, because the literature contains several different notions for this. The collapse result of Chakaravarthy and Roy that we use follows the loose oracle access mode adopted in [CR11]. Namely, oracle queries outside of the promise set are allowed and no particular behaviour of the computational model defining the promise class is expected on such queries. At the same time, the answer of the (base) machine using this oracle must be correct irrespective of the oracle’s answers to such queries and no assumption is made on the internal consistency of the answers outside of the promise set. For deterministic polynomial-time oracle machines this approach is equivalent to querying any language consistent with the promise problem, that is, a language that contains all the “yes” instances and does not contain the “no” instances of the promise problem. One can also extend it to the classes of promise problems: A class $\mathcal{C}$ of languages is consistent with a class $\mathcal{D}$ of promise problems if for every problem $\Pi\in\mathcal{D}$ there is a language $L\in\mathcal{C}$ consistent with $\Pi$.
The equivalence between the approaches follows from the works of [GS88, BF99], where the latter approach has been adopted. Nonetheless, we include a formal proof of this equivalence in Appendix C for the sake of completeness.
@@ 1.3 Our Contribution
1.3.1 A New Lower Bound for $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ and the Strongest Non–Input-Oblivious Karp–Lipton Collapse
The containment $\mathbf{{P}}^{\mathbf{{prMA}}}\subseteq\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ and its consequences.
@@
  • whether a Karp–Lipton–style theorem holds for $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ [KP24],
  • whether $\mathbf{{P}}^{\mathbf{{prMA}}}$ is contained in $\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ [CR11].
@@
Combining this theorem with a result of Chakaravarthy and Roy [CR11] that $\mathbf{{NP}}\subseteq\mathbf{{P/poly}}$ implies the collapse $\mathbf{{PH}}=\mathbf{{P}}^{\mathbf{{prMA}}}$, we obtain a Karp–Lipton–style collapse theorem for $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$, thus solving an open question posed in [KP24].
@@
Which Karp–Lipton–style collapse is stronger?
Theorem 1 shows that $\mathbf{{P}}^{\mathbf{{prMA}}}$ is currently the smallest non–input-oblivious class for which a Karp–Lipton–style collapse is known. On the other hand, such a collapse is known for $\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ [CR06], which is input-oblivious. However, the exact relationship between $\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ and $\mathbf{{P}}^{\mathbf{{prMA}}}$ remains unknown. We demonstrate how to unify these results by observing that the collapse can be extended to $\mathbf{{P}}^{\mathbf{{prOMA}}}\subseteq\mathbf{{P}}^{\mathbf{{prMA}}}$, where $\mathbf{{prOMA}}$ is the input-oblivious version of $\mathbf{{prMA}}$ and therefore is contained in $\mathbf{{prMA}}$. At the same time, we show that $\mathbf{{P}}^{\mathbf{{prOMA}}}$ is also contained in $\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$, thus making $\mathbf{{P}}^{\mathbf{{prOMA}}}$ smaller than both $\mathbf{{P}}^{\mathbf{{prMA}}}$ and $\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$! As an intermediate step we show that $\mathbf{{P}}^{\mathbf{{prO_{2}^{\mathbf{{\scriptscriptstyle P}}}}}}\subseteq\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$.
Theorem 3.
$\mathbf{{P}}^{\mathbf{{prOMA}}}\subseteq\mathbf{{P}}^{\mathbf{{prO_{2}^{\mathbf{{\scriptscriptstyle P}}}}}}\subseteq\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$.
We note that the “non-promise” version of the latter inclusion, i.e. $\mathbf{{P}}^{\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}}\subseteq\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$, was already established in [CR06]. Yet, this result does not carry over to the promise case. A similar phenomenon occurs with the non–input-oblivious analogue of this question: while we know that $\mathbf{{P}}^{\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}}\subseteq\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ [CR06], it still remains open whether $\mathbf{{P}}^{\mathbf{{prS_{2}^{\mathbf{{\scriptscriptstyle P}}}}}}\subseteq\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$.

### v2

@@ Abstract
Korten and Pitassi (FOCS, 2024) defined a new [^1] Note that this notation had been used in the past [Sch83] for a very different class, which has been apparently forgotten after that. complexity class $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ as the polynomial-time Turing closure of the Linear Ordering Principle. They put it between $\mathbf{{MA}}$ (Merlin–Arthur protocols) and $\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ (the second symmetric level of the polynomial hierarchy).
In this paper we sandwich $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ between $\mathbf{{P}}^{\mathbf{{prMA}}}$ and $\mathbf{{P}}^{\mathbf{{prSBP}}}$. (The oracles here are promise problems, and $\mathbf{{SBP}}$ is the only known class between $\mathbf{{MA}}$ and $\mathbf{{AM}}$.) The containment in $\mathbf{{P}}^{\mathbf{{prSBP}}}$ is proved via an iterative process that uses a $\mathbf{{prSBP}}$ oracle to estimate the average order rank of a subset and find the minimum of a linear order.
Another containment result of this paper is $\mathbf{{P}}^{\mathbf{{prO_{2}^{\mathbf{{\scriptscriptstyle P}}}}}}\subseteq\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ (where $\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ is the input-oblivious version of $\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$). These containment results altogether have several byproducts:
  • We give an affirmative answer to an open question posed by Chakaravarthy and Roy (Computational Complexity, 2011) whether $\mathbf{{P}}^{\mathbf{{prMA}}}\subseteq\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$, thereby settling the relative standing of the existing (non-oblivious) Karp–Lipton–style collapse results of [CR11] and [Cai07],
  • We give an affirmative answer to an open question of Korten and Pitassi whether a Karp–Lipton–style collapse can be proven for $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$,
  • We show that the Karp–Lipton–style collapse to $\mathbf{{P}}^{\mathbf{{prOMA}}}$ is actually better than both known collapses to $\mathbf{{P}}^{\mathbf{{prMA}}}$ due to Chakaravarthy and Roy (Computational Complexity, 2011) and to $\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ also due to Chakaravarthy and Roy (STACS, 2006). Thus we resolve the controversy between previously incomparable Karp–Lipton collapses stemming from these two lines of research.
@@ 1 Introduction
The seminal theorem of Richard M. Karp and Richard J. Lipton [KL80] connected non-uniform and uniform complexity by demonstrating a collapse of the Polynomial Hierarchy assuming $\mathbf{{NP}}$ has polynomial-size Boolean circuits. This collapse has since been very instrumental in transferring lower bounds against Boolean circuits of fixed-polynomial [^2] That is, for any $k\in\mathbb{N}$, the class contains a language that cannot be computed by Boolean circuits of size $n^{k}$, i.e. a language outside of $\mathsf{Size}[n^{k}]$. size to smaller classes of the Polynomial Hierarchy. Since then, these results were strengthened in many ways leading to “minimial” complexity classes that have such lower bounds and to which the Polynomial Hierarchy collapses.
@@ 1.1.1 Classes Based on Symmetric Alternation
An important notion in this context is that of symmetric alternation. Namely, one of the best collapses was based on the following idea ([Cai07], attributed to Sengupta): if polynomial-size circuits for SAT exist, two provers (defending the answers ‘yes’ and ‘no’, respectively) send such circuits to a polynomial-time bounded verifier who can, in turn, use them to verify membership in any language in $\mathbf{{PH}}$. The corresponding class $\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ [Can96, RS98] was thus shown to have fixed-polynomial circuit lower bounds. (In Section 2 we provide formal definitions for all less known classes we use.)
Indeed, since $\mathbf{{NP}}\subseteq\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$, if SAT requires superpolynomial circuits, we are done. Otherwise, the Polynomial Hierarchy, which is known to contain “hard” languages (that is, for every $k\in\mathbb{N}$, $\mathbf{{PH}}\not\subseteq\mathsf{Size}[n^{k}]$) by Kannan’s theorem [Kan82], collapses to $\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ and so do these hard languages. This technique has been known as a win-win argument in the literature [Kan82, BCG+96, KW98, Vin05, Cai07, San09, CR11, Vol14, IKV23]. Chen et al. [CMMW19] prove that there is a bidirectional relationship between fixed-polynomial lower bounds and Karp–Lipton–style theorems. In the linear-exponential regime, while the win-win argument can be extended to obtain superpolynomial lower bounds for $\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle E}}}}}$ (the linear-exponential version of $\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$), it falls short of achieving truly exponential lower bounds, as it encounters the so-called “half-exponential” barrier (see [MVW99]).
@@
An important feature of these new results was that they were based on reducing finding a hard function to a total search problem. Namely, the works of Korten [Kor22] and Li [Li24] reduced the question to the so-called Range Avoidance problem: given a function $f\colon\{0,1\}^{n}\to\{0,1\}^{m}$ with $m>n$, represented by a Boolean circuit, find a point outside its image. In [CHR24, Li24], Range Avoidance has been reduced to symmetric alternation. Subsequently, Korten and Pitassi [KP24] reduced Range Avoidance to the Linear Ordering Principle: given an implicitly described ordering relation, either find the smallest element or report a breach of the linear order axioms. A polynomial-time Turing closure of this principle gave rise to a new class $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}\subseteq\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$: a version of $\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ where the two provers provide points of a polynomial-time verifiable linear order on binary strings of a certain length (each point starting with the corresponding answer 0 or $1$), and the prover that provides the smaller element wins.
@@ 1.1.2 Classes Based on Merlin-Arthur Protocols
In a parallel line of research, the same questions were considered for classes based on Merlin–Arthur proofs: Santhanam [San09] has shown fixed-polynomial lower bounds for promise problems possessing such proofs (i.e. the class $\mathbf{{prMA}}$). In [CR11], Chakaravarthy and Roy have shown a Karp–Lipton–style collapse and thus fixed-polynomial size lower bounds for the class $\mathbf{{P}}^{\mathbf{{prMA}}}$. In particular, they presented a new upper bound for $\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ by showing that $\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}\subseteq\mathbf{{P}}^{\mathbf{{prAM}}}$. Nonetheless, the relationship between $\mathbf{{P}}^{\mathbf{{prMA}}}$ and the classes of symmetric alternation (including $\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$, $\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$, and then-unknown $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$) remained open.
Combining their upper bound for $\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ with a result of [AKSS95], that $\mathbf{{NP}}\subseteq\mathbf{{P/poly}}$ implies an “internal collapse” $\mathbf{{MA}}=\mathbf{{AM}}$ [^3] The internal collapse goes through for the promise versions of the classes as well., [CR11] concluded that the Polynomial Hierarchy collapses all the way to $\mathbf{{P}}^{\mathbf{{prMA}}}$. Subsequently, by applying the win-win argument, they obtained fixed-polynomial bounds for $\mathbf{{P}}^{\mathbf{{prMA}}}$, which (unlike $\mathbf{{prMA}}$) is a class of languages. It is to be noted though that since $\mathbf{{prMA}}$ is not a class of languages — while $\mathbf{{P}}^{\mathbf{{prMA}}}$ is, there is no immediate way to carry any lower bound against $\mathbf{{prMA}}$ over to $\mathbf{{P}}^{\mathbf{{prMA}}}$: it is not clear how to leverage (even) Turing reductions to construct a specific language consistent with a given promise problem.
Babai, Fortnow, and Lund [BFL91] prove that if $\mathbf{{EXP}}\subseteq\mathbf{{P/poly}}$, then $\mathbf{{EXP}}=\mathbf{{MA}}$. Although this is a much larger class, the proof has the advantage that it does not relativize. More collapses in the exponential regime have been proved since then [IKW02, BH92], and the win-win argument yields superpolynomial lower bounds for some of them: $\mathbf{{MAEXP}}\nsubseteq\mathbf{{P/poly}}$ [BFT98].
@@ 1.2 Promise Problems as Oracles
An important note is due on the use of a promise problem as an oracle, because the literature contains several different notions for this. The collapse result of Chakaravarthy and Roy that we use follows the loose oracle access mode adopted in [CR11]. Namely, oracle queries outside of the promise set are allowed and no particular behaviour of the computational model defining the promise class is expected on such queries. At the same time, the answer of the (base) machine using this oracle must be correct irrespective of the oracle’s answers to such queries and no assumption is made on the internal consistency of the answers outside of the promise set.
For deterministic polynomial-time oracle machines this approach is equivalent to querying any language consistent with the promise problem, that is, a language that contains all the “yes” instances and does not contain the “no” instances of the promise problem. One can also extend it to complexity classes: a class $\mathcal{C}$ of languages is consistent with a class $\mathcal{D}$ of promise problems if for every problem $\Pi\in\mathcal{D}$ there is a language $L\in\mathcal{C}$ consistent with $\Pi$. The equivalence between the approaches follows from the works of [GS88, BF99], where the latter approach has been adopted. Nonetheless, we include a formal proof of this equivalence in Section 2.1 for the sake of completeness and for further reference.
@@ 1.3 Our Contribution
In this paper we prove the inclusions $\mathbf{{P}}^{\mathbf{{prMA}}}\subseteq\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}\subseteq\mathbf{{P}}^{\mathbf{{prSBP}}}$ and $\mathbf{{P}}^{\mathbf{{prO_{2}^{\mathbf{{\scriptscriptstyle P}}}}}}\subseteq\mathbf{{O_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$, which not only give new upper and lower bounds for $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$, but also demonstrate that the Karp–Lipton–collapse to $\mathbf{{P}}^{\mathbf{{prOMA}}}$ is currently the best one both for symmetric–alternation–based and Merlin–Arthur–based classes of languages.
1.3.1 A New Lower Bound for $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ and the Strongest Non-Input-Oblivious Karp–Lipton Collapse
@@
  • whether $\mathbf{{P}}^{\mathbf{{prMA}}}$ is contained in $\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ [CR11] (note that for these two classes Karp–Lipton style theorems have been proved by [CR11] and by Samik Sengupta [unpublished], respectively),
  • whether a Karp–Lipton–style theorem holds for $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$ [KP24].
@@
Combining this theorem with a result of Chakaravarthy and Roy [CR11] that $\mathbf{{NP}}\subseteq\mathbf{{P/poly}}$ implies the collapse $\mathbf{{PH}}=\mathbf{{P}}^{\mathbf{{prMA}}}$, we obtain a Karp–Lipton–style collapse theorem for $\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}$, thus resolving an open question posed in [KP24].
@@
Together with the result of [Cai07], it lines up all known non-input-oblivious classes for which a Karp–Lipton–style collapse has been shown:
$\mathbf{{P}}^{\mathbf{{prMA}}}\subseteq\mathbf{{L_{2}^{\mathbf{{\scriptscriptstyle P}}}}}\subseteq\mathbf{{S_{2}^{\mathbf{{\scriptscriptstyle P}}}}}\subseteq\mathbf{{ZPP}}^{\mathbf{{NP}}}\subseteq\mathbf{{\Sigma_{2}^{\mathbf{{\scriptscriptstyle P}}}}}.$